OpenStudy (anonymous):

when n->infinit: limit [1/(n^2)+1/(n+1)^2+...+1/(2n)^2]=? My answer is 0, but the textbook is 1. Am I right?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{1}{n^2} + \frac{1}{n^2+1}+...+\frac{1}{2n^2}\]

OpenStudy (anonymous):

No the second item is 1/[(n+1)^2]

OpenStudy (anonymous):

Oh Okay

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{1}{n^2} + \frac{1}{(n+1)^2}+...+\frac{1}{(2n)^2}\]

OpenStudy (anonymous):

It is zero ...right?

OpenStudy (anonymous):

Looks zero to me too.

OpenStudy (anonymous):

wait satellite is here

OpenStudy (anonymous):

not sure it is zero

OpenStudy (anonymous):

Thank you for your confirm..

OpenStudy (anonymous):

but it is not obvious to me what it is. requires some work yes?

OpenStudy (anonymous):

not sure it is zero

OpenStudy (anonymous):

I use squeez theorem let Sn=this series then: 1/4n = [1/(2n)^2+1/(2n)^2+...1/(2n)^2]< Sn <[1/n^2+1/n^2+...1/n^2]=1/n

OpenStudy (anonymous):

then n-> infinit 0<Sn<0. SO Sn=0 ? But the Squeeze Theorem need 0<=Sn<=0 ... SO I am not sure

OpenStudy (anonymous):

Or just compute the limit for each item, then 0+0+0...=0

OpenStudy (anonymous):

yeah looks good to me!

OpenStudy (anonymous):

no your first method looks good. you have \[\frac{1}{n^2}+...+\frac{1}{(2n)^2}<\frac{1}{n^2}+...+\frac{1}{n^2}=\frac{1}{n}\]

OpenStudy (anonymous):

be careful adding infinite number of things that tend to zero. not really a method. i like your answer above

OpenStudy (anonymous):

you type the equation by code. Is there any help so I can learn it ?

OpenStudy (anonymous):

Yes, thank you for your reminding "be careful adding infinite number of things that tend to zero".

OpenStudy (anonymous):

Use \ [ without the space to start an equation. \ ] is used to close an equation. Within the equation \frac{}{} is used to display a fraction. ^{} is used for exponents.